OpenStudy (anonymous):

An object spring system moving with simple harmonic motion has an amplitude A. a) What is the total energy of the system in terms of k and A only? (E=1/2kA^2) b) Suppose at a certain instant the kinetic energy is twice the elastic potential energy. Write an equation describing this situation, using only the variables for the mass m, velocity v, spring constant k, and position x. (1/2mv^2=kx^2) c) Using the results of parts (a) and (b) and the conservation of energy equation, find the positions x of the object when its kinetic energy equals twice the potential energy stored in the spring.

OpenStudy (anonymous):

I only need part c. The answers to parts a and b are shown in parenthesis.

OpenStudy (abb0t):

\[T = 1.5 U\] \[\frac{1}{2}K (A^2-x^2) = 1.5 \frac{1}{2}Kx^2\] \[A^2-x^2 = 1.5x^2\] \[2.5x^2=A^2\] Can you get position now?

OpenStudy (anonymous):

Ah yes. Thanks!

OpenStudy (anonymous):

quick question...what does the U stand for...

OpenStudy (anonymous):

@abb0t